the sum of two numbers is 10. the first number is 2/3 the second. what are the numbers?
Answer by ptth176(50)

the first number=x
the second number=y
x+y=10(1)
x=2/3y(2)
(1)&(2)=>2/3y+y=10
=>5/3y=10
=>y=6
=>x=4
Finally,the first number=4
the second number=6